OpenStudy (anonymous):

A baseball team has home games on Wednesday and Sunday. Then two games together earn $5110.00 for the team. Wednesday's games generate $490.00 less then Sunday's game. How much money was taken in each game?

OpenStudy (anonymous):

someone help me please!

OpenStudy (anonymous):

do you understand the question?

OpenStudy (nurali):

Let S, W be the earmings on Sunday, Wednesday. S+W=5110.00 W=S-490.5. S+S-490.5=5110.00 2S=5110.00+490.5 2S=5600.5 S=5600.5/2 =2800.25 W=2800.25-490.5 =$2309.75

OpenStudy (nurali):

Welcome to Openstudy.

OpenStudy (anonymous):

thank you so much!

OpenStudy (nurali):

My Pleasure.